### v2

@@ 4.2 Comparison to the State of the Art
Table 1: Performance comparison with the state-of-the-art models GDanceR [30] and GCD [29]. Leveraging the advancements in single-dancer modeling from EDGE (EDGE), we adapted it into a multi-dancer model by training it on the AIOZ-GDance dataset. Our approach attained superior results on the Group-dance Metric, concurrently mitigating foot slip occurrences.
@@
EDGE [55] 63.35 61.72 0.36 31.40 9.57 0.26 2.63
GDanceR [30] 51.27 79.01 0.22 43.90 9.23 0.25 3.05
GCD [29] 31.47 80.97 0.17 31.16 10.87 0.26 2.53
TCDiff (Ours) 21.12 81.48 0.15 42.39 14.37 0.25 0.54
@@
GDanceR[30] 2 53.83 75.44 0.286 48.82 9.36 0.248
3 55.85 74.07 0.204 44.47 9.36 0.245
4 58.79 77.71 0.162 47.32 9.24 0.248
5 55.05 78.72 0.218 44.19 8.99 0.249
GCD[29] 2 34.09 80.26 0.167 32.62 10.41 0.266
3 36.25 79.93 0.184 33.94 10.02 0.266
4 36.28 81.82 0.125 35.89 9.87 0.251
5 38.43 81.44 0.168 35.08 9.92 0.264
TCDiff (Ours) 2 44.69 78.94 0.190 60.93 13.30 0.247
3 10.97 80.51 0.123 48.00 19.28 0.253
4 13.44 80.70 0.149 23.32 10.89 0.253
5 15.36 85.77 0.109 37.31 14.01 0.236

### v3

@@ 4.2 Comparison to the State of the Art
Table 1: Performance comparison with the state-of-the-art models GDanceR [30], GCD [29] and CoDancers [60]. Leveraging the advancements in single-dancer modeling from EDGE[55], we adapted it into a multi-dancer model by training it on the AIOZ-GDance dataset. Our approach attained superior results on the Group-dance Metric, concurrently mitigating foot slip occurrences. Red and Blue indicate the best and the second best result.
@@
EDGE [55] 70.26 62.24 0.42 33.24 10.11 0.26 3.02
GDanceR [30] 46.01 77.22 0.20 49.37 9.34 0.25 3.05
GCD [29] 31.95 80.55 0.16 37.70 9.64 0.26 2.53
CoDancers [60] 26.10 74.05 0.10 23.98 9.48 0.25 3.26
TCDiff (Ours) 13.86 81.98 0.13 37.47 15.10 0.25 0.51
@@
GDanceR[30] 2 48.77 78.23 0.203 50.84 9.01 0.248
3 45.88 75.61 0.224 49.69 9.75 0.247
4 44.63 76.83 0.161 47.19 9.54 0.246
5 44.76 75.22 0.198 49.75 9.04 0.247
GCD[29] 2 30.10 79.36 0.125 37.28 9.78 0.260
3 33.26 79.78 0.163 36.24 9.87 0.260
4 34.23 82.83 0.172 38.07 9.45 0.251
5 30.22 80.22 0.185 39.24 9.45 0.260
CoDancers[60] 2 24.53 72.88 0.080 26.31 9.01 0.251
3 27.23 74.34 0.084 24.85 9.15 0.254
4 26.44 75.34 0.097 25.76 9.43 0.258
5 26.34 74.22 0.113 25.45 9.77 0.253
TCDiff (Ours) 2 15.77 81.92 0.121 41.26 16.20 0.263
3 10.97 81.51 0.123 48.00 19.28 0.253
4 13.44 81.70 0.149 23.32 10.89 0.253
5 15.36 82.77 0.109 37.31 14.01 0.236